# Dynamic linear regression models for forecasting time series with semi long memory errors

### v1

@@ 5.3. Forecasting evaluation settings
Given the abundance of training data, i.e. a very large $T$, the posterior distribution $p(\boldsymbol{\theta}|y_{1:T-h-i})$ remains fairly unchanged for varying values $h\leq 15$ and $i\leq k=100$. Thus, for computational convenience, we avoid reestimating the model and sample $10{,}000$ MCMC draws from the posterior conditional on the full training data. The same posterior samples are then used to compute (5.2), i.e. only the conditional likelihood part of the integrand changes. The distributional forecasts are assessed via the continuous rank probability score (CRPS), which is defined as
@@
where $F$ is the cumulative distribution function of the forecast and $\mathbbm{1}$ is the indicator function, taking the value one if the argument is true or zero otherwise (Matheson and Winkler,, 1976). When the distributional forecasts $F$ and the data distribution are equal, the CRPS achieves its minimum. We use 300 samples from the posterior predictive distribution in (5.2) to estimate the empirical cumulative distribution function $\widehat{F}(\widetilde{y}_{T+i})$ for all, $i=1,\dots,k=100$. The estimated CRPS is then summed over all time points to obtain $\widehat{\text{CRPS}}^{(h)}$, for $h=1,\dots,15$.
@@ 5.4. Example 1: New England electricity demand
The New England data consists of hourly electricity demand and temperature data for the states of Maine and Vermont from 1st March 2003 to April 30th, 2017, with $T=124{,}200$ observations. The response, Maine electricity demand, $y_{t}$, is modelled as a linear combination of lagged exogenous variables, Maine temperature $x_{1,t-1}$ and Vermont (neighbouring state) electricity demand $x_{2,t-1}$. Strong multi-seasonal and trend components exist within each variable. The mstl function in the R package Forecast (Hyndman and Khandakar,, 2008) is used to remove these effects. Figure 5 depicts each variable after removing the multi-seasonal and trend components and applying a Box-Cox transformation (Box and Cox,, 1964). As seen from the figure, the response exhibits long memory, which makes it a suitable candidate for an ARFIMA or ARTFIMA error process.
Figure 5. Maine and Vermont electricity demand (megawatts) and temperature (degrees Celsius) time series data after removing multi-seasonal and trend components and applying a Box-Cox transformation, alongside their corresponding autocorrelation plots. The measurement units are for the raw data.
The search over the model space described in Section 5.2 resulted in the following lag-orders for each error process: ARTFIMA$(1,d,\lambda,2)$, ARFIMA$(3,d,1)$, ARMA$(3,1)$. We thus have two cases: $p=1,q=2$ and $p=3,q=1$.
Table 2 reports the BIC values and average computation time for a log-likelihood evaluation for all three error processes for the case $p=1,q=2$. The table shows that ARFIMA has the smallest BIC value, however, it is roughly 40 times slower than both ARFIMA and ARMA since it uses the Gaussian log-likelihood. ARTFIMA has the second smallest BIC value and is only marginally slower than the ARMA model.
${\text{ARTFIMA}(1,d,\lambda,2)}$ ${\displaystyle\text{ARFIMA}(1,d,2)}$ ${\displaystyle\text{ARMA}(1,2)}$
BIC -383273.75 -383578.45 -382749.34
@@
Table 2. BIC values and the average time for one log-likelihood evaluation for each dynamic linear regression model for New England electricity demand.
@@
Figure 6 plots the RMSE and CRPS for each error process for different forecast horizons $h$. The RMSE for all models are similar for small $h$; however, $h>6$ shows a lower RMSE score for the ARMA and ARTFIMA models compared to ARFIMA, with the ARMA model performing the best for $h>7$. The CRPS is similar for all models, with the ARFIMA model having the overall lowest and ARMA performing worse over the forecast horizon. Figure 7 plots the predictions and prediction intervals for the three error processes when predicting the electricity demand one and six hours ahead, i.e. $h=1,6$. All three error processes perform accurate predictions, with the prediction intervals getting wider for larger horizons as expected. To summarise the results, we conclude that ARMA and ARTFIMA are computationally much more efficient than ARFIMA. ARMA performs surprisingly well for forecasting, which indicates that the long-range dependency feature of the ARTFIMA and ARFIMA models does not strongly influence the forecasts in this case. However, inspecting the posterior density of $\beta$ for the different error processes in Figure 8 reveals a significant difference.
@@
Figure 6. Root mean square error and the continuous rank probability scores for $h$-step ahead forecasts for New England electricity demand for $p=1,q=2$.
@@
Figure 7. Prediction and prediction intervals for one hour ahead ($h=1$, top panel) and ($h=6$, bottom panel) predictions for three error processes (columns). The results are shown for New England electricity demand for $p=1,q=2$.
@@
Figure 8. Kernel density estimator of the posterior density of $\boldsymbol{\beta}$. using $10{,}000$ MCMC samples. The results are shown for the New England data with $p=1,q=2$ and three error processes, ARTFIMA, ARFIMA, and ARMA.
Next, we consider the case of $p=3,q=1$. The MCMC sampler had problems converging for the ARTFIMA model; thus, the CRPS was computed at the MAP instead of using the full posterior predictive distribution. Table 3 displays the BIC values for each model (the relative computing times are similar to the $p=1$ $q=2$ case and therefore omitted). ARTFIMA has the smallest BIC value, followed by ARFIMA and ARMA. However, in terms of RMSE forecasts in Figure 9, ARMA performs the best. In terms of CRPS, both ARTFIMA and ARFIMA perform better than ARMA.
${\text{ARTFIMA}(3,d,\lambda,1)}$ ${\displaystyle\text{ARFIMA}(3,d,1)}$ ${\displaystyle\text{ARMA}(3,1)}$
BIC -383696.40 -383608.18 -382959.42
Table 3. BIC values for each dynamic regression model for New England electricity demand.
Figure 9. Root mean square error and the continuous rank probability scores for $h$-step ahead forecasts for New England electricity demand for $p=3,q=1$.
@@
Figure 10 plots the fitted spectral densities (log-log scale) of the dynamic regression models with ARTFIMA and ARFIMA errors together with their corresponding periodogram data. Due to the dependence on $\boldsymbol{\beta}$, the two periodograms are different for every frequency $\omega$, but are almost identical at the lower frequencies. Consistent with the results of Sabzikar et al., (2019), the ARTFIMA spectrum provides a better overall fit to the periodogram data, in particular for small frequencies.
@@
Figure 10. Spectral densities at the MAP and their respective periodograms. The results are shown for the New England data with $p=3,q=1$. The ARTFIMA spectrum (black line) with its periodogram (grey circles) and the fitted ARFIMA (orange line) with its corresponding periodogram (orange dots).
@@ 5.5. Example 2: Victorian electricity demand
The second application we consider is the half-hourly electricity demand for Victoria, Australia. The data consists of $T=52{,}608$ observations of electricity demand (in megawatts) and Melbourne’s temperature (degrees Celsius) between the 1st January 2012 - 31st December 2014. The response $y_{t}$, Victorian electricity demand, is modelled as a linear combination of the lagged exogenous variable Melbourne temperature $x_{t-1}$. Strong multi-seasonal and trend components exist within each variable and were removed via the mstl function. Figure 11 depicts each variable after removing the multi-seasonal and trend components and applying a Box-Cox transformation.
@@
Figure 11. Victoria, Australia electricity demand (in megawatts) and temperature (degrees Celsius) time series data after removing multi-seasonal and trend components and applying a Box-Cox transformation, alongside their corresponding autocorrelation plots. The measurement units are for the raw data.
Despite removing seasonality, as seen from Figure 11, all seasonality cannot be removed and still exists at multiples of the 48th lag (1 day is 48 half-hours). This is also present in the periodogram, see Figure 14, which peaks at the corresponding frequencies.
All models achieve their smallest BIC for $p=2,q=1$ following the procedure described in Section 5.2. In addition, we include one seasonal MA term $Q=1$ with $s=48$ for ARMA and ARTFIMA and denote this parameter $\Psi^{\star}_{1}$ as in (2.3). We do not include a seasonal ARFIMA model as the Gaussian likelihood is expensive to compute and our paper has overwhelming evidence that ARTFIMA captures sufficient long-range dependence. Table 4 displays the BIC and the average time for one log-likelihood evaluation for the five models. The seasonal models have substantially lower BIC values than the non-seasonal models. The seasonal ARTFIMA error process gives the smallest BIC, followed by the seasonal ARMA. The table also shows that significant computational speedups are gained for ARTFIMA and ARMA models due to the frequency domain approach.
${\scriptstyle\text{ARTFIMA}(2,d,\lambda,1)(0,0,1)_{48}}$ $\scriptstyle{\text{ARTFIMA}(2,d,\lambda,1)}$ ${\scriptstyle\text{ARFIMA}(2,d,1)}$ ${\scriptstyle\text{ARMA}(2,1)}$ ${\scriptstyle\text{ARMA}(2,1)(0,0,1)_{48}}$
BIC -525257.64 -512383.65 -512102.13 -511832.00 -520631.12
Time 35.1ms 3.96ms 203.0ms 3.79ms 35.1ms
@@
Table 4. BIC values for each dynamic regression model for Victorian electricity demand.
This dataset is particularly challenging and the MAP estimates were on the border of the stationarity or invertibility region (when transformed back from the partial autocorrelation parameterisation) for all models. As a consequence, MCMC sampling was difficult to carry out with reasonable efficiency. Thus, the CRPS (that requires posterior sampling) was only computed using the MAP estimate. Figure 12 reports the forecast RMSE and CRPS for the five models for all horizons $h=1,\dots,15$. The seasonal models have the lowest RMSE and CRPS compared to the non-seasonal models. The seasonal ARTFIMA model has the best forecasting ability in terms of both point estimation (RMSE) and distributional forecasts (CRPS). Figure 13 plots the predictions and prediction intervals for the seasonal models when predicting the electricity demand one and six hours ahead, i.e. $h=2,12$. While the one hour ahead prediction is accurate for both models (albeit slightly more accurate for ARTFIMA), the seasonal ARTFIMA clearly performs better for the six hours ahead forecast.
@@
Figure 12. Root mean square error and the continuous rank probability score of the $h$-step ahead forecasts for Victorian electricity data for $p=2,q=1$
@@
Figure 13. Prediction and prediction intervals for one hour ahead ($h=2$, top panel) and ($h=12$, bottom panel) predictions for two seasonal error processes (columns). The results are shown for Victorian electricity data for $p=2,q=1$.
@@
Finally, Figure 14 further shows the appropriateness of the seasonal ARTFIMA model by displaying a log-log plot of the periodogram and spectral density at the MAP. It is evident that the seasonal ARTFIMA model fits the lower frequencies $\text{log}(\omega)<-4$ well, and also captures the seasonal features of the periodogram.
@@
Figure 14. Main plot: Log frequency vs log power (log-log) plot of the periodogram and the spectral density function at the MAP for DLR with ARTFIMA errors. Zoomed-in plot: Similar figure but displaying moderate frequencies on a linear frequency scale.
@@ 6. Conclusion and future research
We proposed a dynamic linear regression models with an ARTFIMA error process that accounts for long-range dependence. We developed a frequency domain estimation method that utilises the computationally efficient FFT. The frequency domain method uses a likelihood based on asymptotic properties of the periodogram, and we demonstrated empirically that the resulting inference matches well with time domain likelihood methods. This result was expected as the case without an exogenous variable has strong theoretical results regarding the approximation error vanishing asymptotically.
We showed empirically that the well-known phenomena of the spectral density of the ARFIMA model not fitting the periodogram data well for small frequencies carries over to the dynamic linear regression case. As such, we argue that dynamic linear regression models with ARFIMA errors are not well suited for frequency domain methods, and their estimation needs time domain methods which are computationally expensive. In contrast, we showed that the dynamic linear regression models with ARTFIMA errors fit the low-frequency spectrum well and are thus a computationally fast alternative to model long-range dependence.
Our proposed model demonstrated superior forecasting performance when the data exhibited complex long-range dependency and was marginally worse if not on par with the ARMA error process for less complex data. This is expected as our model nests the dynamic linear regression model with ARMA errors. We did not see any signs of overfitting when using an ARTFIMA error process instead of an ARMA for the simpler dataset. In addition, we also observed that the posterior distribution of the regression parameter could be significantly affected by not accounting for long-range dependency in the error process. In summary, we believe dynamic linear regression models with ARTFIMA errors are robust methods that should be routinely included in a forecaster’s toolkit.
Future research will extend our approach to modelling a multivariate response as in Villani et al., (2022). For small to moderate datasets, the Whittle log-likelihood is known to be severely biased. Exploring recent advances such as the debiased Whittle likelihood (Sykulski et al.,, 2019) for dynamic linear regression models can potentially enable our method for small to moderately large datasets. Finally, the structure of our model can utilise subsampling MCMC methods (Quiroz et al.,, 2019, 2021) in the spectral domain (Salomone et al.,, 2020).

### v2

@@ 5.3. Settings for forecast evaluation
$\approx\log\left(\frac{1}{M}\sum^{M}_{m=1}p(y_{T+h+i}|\boldsymbol{\theta}^{(m)},y_{1+i:T+i})\right),$
and the last line is a Monte Carlo approximation thereof, with $\boldsymbol{\theta}^{(m)}\sim p(\boldsymbol{\theta}|y_{1+i:T+i})$ and $M=900$. The LPDS gives a measure of how well the model fits the out-of-sample data (Gelman et al.,, 2014), with larger values being preferred. The other two measures (RMSE and CRPS), however, have a reciprocal scale (smaller is better), and thus, we report the negative LPDS so that smaller values are preferred for all three metrics.
To assess point forecasts from a given model, define the conditional expectation when predicting the $i$th out-of-sample observation as
@@
where $p(\widetilde{y}_{T+h+i}|y_{1+i:T+i})$ is the posterior predictive distribution. To evaluate the performance of the point forecasts, the root mean square error (RMSE) is computed as
$\text{RMSE}^{(h)}=\sqrt{\frac{1}{k-h+1}\sum^{k-h}_{i=0}\left(y_{T+h+i}-\widehat{y}_{T+h+i}\right)^{2}},$
for each forecast horizon $h$. The RMSE is the square root of the average squared distance between the out-of-sample observations and their forecasted values, and thus, smaller values are preferred.
The distributional forecasts $F()$ are assessed via the continuous rank probability score (CRPS), which is defined as
[…]
where $\mathbf{1}$ is the indicator function, taking value one if $y_{T+h+i}\leq\widetilde{y}_{T+h+i}$ or zero otherwise (Matheson and Winkler,, 1976). When the distributional forecasts $F$ and the data distribution are equal, the CRPS achieves its minimum and thus smaller values are preferred. The distributional forecasts are unknown but can be estimated using samples from the $h$-step-ahead predictive posterior distribution to estimate the empirical cumulative distribution function $\widehat{F}(\widehat{y}_{T+h+i})$ for all, $i=0,\dots,k-h$. The final CRPS is formed by taking the mean over the out-of-sample observations, i.e.
$\text{CRPS}^{(h)}=\frac{1}{k-h+1}\sum^{k-h}_{i=0}\text{CRPS}^{(h)}_{i}(F,y_{T+h+i}).$
To compute the RMSE and CRPS, we use 900 samples from the posterior predictive.
@@ 5.4. Example 1: New England electricity demand
The New England data consists of hourly electricity demand (megawatts) and temperature (degrees Celsius) for the states of Maine and Vermont from 1st March 2003 to April 30th, 2017, with $T=124{,}200$ observations presented in Hong et al., (2016). The response, Maine electricity demand $y_{t}$, is modelled as a linear combination of lagged exogenous variables, Maine temperature $x_{1,t-1}$ and Vermont (neighbouring state) electricity demand $x_{2,t-1}$. Strong multi-seasonal and trend components exist within each variable. The mstl function in the R package Forecast (Hyndman and Khandakar,, 2008) is used to remove these effects. Figure 5 plots each variable after removing the multi-seasonal and trend components and applying a Box-Cox transformation (Box and Cox,, 1964). The figure shows that the response exhibits some long memory, making it a suitable candidate for an ARFIMA or ARTFIMA error process.
Figure 5. The left panels plot the Maine and Vermont hourly electricity demand and temperature data each after removing multi-seasonal and trend components and applying a Box-Cox transformation. The right panels are the corresponding autocorrelation plots.
The lowest DIC for ARMA and ARTFIMA error processes occurred for $p=2$ and $q=1$ in both cases. Table 3 reports the DIC values for the three models as well as the average time for one evaluation of the log-likelihood function. The table shows that ARTFIMA has the lowest DIC value, followed by ARFIMA, with ARMA having the highest (recall that lower is better). However, ARFIMA is the most computationally demanding model and is roughly 40 times slower than the other two models since the Whittle likelihood cannot be used. ARTFIMA and ARMA have comparable computational time per iteration of the log-likelihood of 11.5ms and 10.5ms, respectively.
ARTFIMA$(2,d,\lambda,1)$ ARFIMA$(2,d,1)$ ARMA$(2,1)$
DIC $\boldsymbol{71987.56}$ 72155.15 73030.91
@@
Table 3. DIC values and the average time (milliseconds) for one log-likelihood evaluation for each dynamic regression model for New England electricity demand. The lowest DIC value is in boldface.
@@
Figure 6 plots the negative LPDS, RMSE and CRPS for each model over the forecast horizon $h$. For the negative LPDS, ARFIMA performs the best here for $h>3$, and the results for ARMA and ARTFIMA are almost identical. Hence, there is no visual distinction between the two. Looking at the RMSE, the ARFIMA model has the lowest score for small $h$; however, when $h>6$, the ARMA and ARTFIMA models have a lower RMSE score compared to ARFIMA. The CRPS is similar for all models, with no model clearly outperforming the others for several forecast horizons. Figure 7 plots the predictions and prediction intervals for the three error processes when predicting the electricity demand one and six hours ahead, i.e. $h=1,6$. All three error processes perform accurate predictions, with the prediction intervals getting wider for larger horizons as expected. To summarise the results, we conclude that ARMA and ARTFIMA overall provide better scores and are computationally much more efficient than ARFIMA. ARMA performs almost identically to ARTFIMA, which indicates that the (semi) long-range dependency feature of ARTFIMA does not strongly influence the forecasts in this case. However, inspecting the posterior density of the regression parameter $\boldsymbol{\beta}$ in Figure 8 reveals a significant difference.
@@
Figure 6. New England electricity data: negative log-predictive density score (LPDS), root mean square error (RMSE) and the continuous rank probability score (CRPS) for all models based on $h$-step ahead forecasts for $p=2,q=1$.
@@
Figure 7. Prediction and prediction intervals for one hour ahead ($h=1$, top panel) and ($h=6$, bottom panel) predictions for three error processes (columns). The results are shown for New England electricity demand for $p=2,q=1$ after adding back the season and trend.
@@
Figure 8. Kernel density estimates of the posterior densities of the regression parameter $\beta$ using $10,000$ MCMC samples. The results are for the New England data with $p=2,q=1$ and two error processes, ARTFIMA and ARMA.
@@
Finally, Figure 9 plots the fitted spectral densities (log-log scale) of the dynamic regression models with ARTFIMA and ARFIMA errors together with their corresponding periodogram data. Due to the dependence on $\boldsymbol{\beta}$ (estimated $\widehat{\boldsymbol{\beta}}$ different for the two processes), the two periodograms are different for every frequency $\omega$ but are almost identical at the lower frequencies. Consistent with the results of Sabzikar et al., (2019), the ARTFIMA spectrum provides a better overall fit to the periodogram data, in particular for small frequencies.
@@
Figure 9. Spectral densities at the MAP and their respective periodograms. The spectrum of DLR with ARTFIMA$(2,d,\lambda,1)$ errors with its periodogram (grey circles) and the DLR with ARFIMA$(2,d,1)$ errors with its corresponding periodogram (orange dots).
@@ 5.5. Example 2: Victorian electricity demand
The second application concerns the half-hourly electricity demand for Victoria, Australia, presented in Hyndman and Athanasopoulos, (2021). The data consists of $T=52{,}608$ observations of electricity demand (megawatts) and Melbourne’s temperature (degrees Celsius) between the 1st January 2012 - 31st December 2014. The response $y_{t}$, Victorian electricity demand, is modelled as a linear combination of the lagged exogenous variable Melbourne temperature $x_{t-1}$. Strong multi-seasonal and trend components exist within each variable and are removed via the mstl function. Figure 10 depicts each variable after removing the multi-seasonal and trend components and applying a Box-Cox transformation.
@@
Figure 10. The left panels plot half-hourly electricity demand and temperature for Victoria, Australia after removing multi-seasonal and trend components and applying a Box-Cox transformation. The right panels are the corresponding autocorrelation plots.
Despite using mstl to remove seasonality, Figure 10 shows that all seasonality has not been removed and still exists at multiples of the 48th lag (1 day is 48 half-hours). This is also present in the periodogram, see Figure 14, which peaks at the corresponding frequencies. For this reason, we include one seasonal MA term $Q=1$ with $s=48$ for ARMA and ARTFIMA and denote this parameter $\Psi^{\star}_{1}$ as in (2.3). We do not include a seasonal ARFIMA model as the Gaussian likelihood is expensive to compute and our paper has overwhelming evidence that ARTFIMA captures sufficient long-range dependence while being computationally much more efficient.
The lowest DIC for the error processes occurred for two sets of lags, $p=2,q=0$ (ARTFIMA) and $p=2,q=1$ (ARMA), which we denote as ARTFIMA$(2,d,\lambda,0)$ and ARMA$(2,1)$. Table 4 displays the DIC value and the average time for one log-likelihood evaluation for all models. For $p=2,q=0$, the lowest DIC (in bold) is the SARTFIMA model, and the worst (highest) is ARMA. ARTFIMA and ARFIMA have similar values for DIC. For the $p=2,q=1$ case, SARTFIMA obtained the lowest DIC (in bold), followed by SARMA. Again, ARTFIMA and ARFIMA produced similar DIC values. Referring to the computation times presented in Table 4, significant computational speedups are gained for ARTFIMA and ARMA models, which had the shortest run time out of all models due to estimation via the Whittle likelihood. This is followed by the seasonal models, which are roughly 10 times slower than their non-seasonal counterparts. The slowest run times are observed for the ARFIMA model via the exact Gaussian log-likelihood, which is approximately 50 times slower than the non-seasonal ARMA and ARTFIMA models.
SARTFIMA ARTFIMA ARFIMA SARMA ARMA
$p=2,q=0$ $\boldsymbol{-341016.47}$ $-328624.02$ $-328612.70$ $-338582.68$ $-327041.24$
$p=2,q=1$ $\boldsymbol{-341541.07}$ $-328514.60$ $-328587.06$ $-338646.40$ $-328085.01$
Time (ms) 35.1ms 3.96ms 203.0ms 35.1ms 3.79ms
@@
Table 4. DIC values and the average time (milliseconds) for one log-likelihood evaluation for each dynamic regression model for Victorian electricity demand.
@@
Figure 11. Victorian electricity data: negative log-predictive density score (LPDS), root mean square error (RMSE) and the continuous rank probability score (CRPS) for all models based on $h$-step ahead forecasts for $p=2,q=0$.
Figure 11 reports the forecast negative LPDS, RMSE and CRPS of the five models for the $p=2,q=0$ case for all horizons $h=1,\dots,15$. The seasonal models are better across all three metrics compared to the non-seasonal models. The seasonal ARTFIMA model has the best forecasting ability for all three metrics, closely followed by SARMA. In contrast, the ARMA model has the highest values (worst) for all three metrics. The ARTFIMA and ARFIMA models were virtually identical, having lower negative LPDS, RMSE and CRPS scores compared to ARMA.
Figure 12. Victorian electricity data: negative log-predictive density score (LPDS), root mean square error (RMSE) and the continuous rank probability score (CRPS) for all models based on $h$-step ahead forecasts for $p=2,q=1$.
For the $p=2,q=1$ case, Figure 12 reports the forecast negative LPDS, RMSE and CRPS of the five models for all horizons $h=1,\dots,15$. The superiority of the SARTFIMA model is more pronounced compared to the $p=2,q=0$ case (Figure 11) across all metrics.
Figure 13 plots the predictions and prediction intervals for the seasonal models SARTFIMA and SARMA when predicting the electricity demand one and six hours ahead, i.e. $h=2,12$. While the one-hour ahead prediction is accurate for both models with seasonal ARTFIMA slightly more accurate, the seasonal ARTFIMA clearly performs better for the six-hours ahead forecast.
@@
Figure 13. Prediction and prediction intervals for one hour ahead ($h=2$, top panels) and six hours ahead ($h=12$, bottom panels) predictions for the two seasonal error processes (columns) for $p=2,q=1$. The results are shown for Victorian electricity data after adding back the trend and seasonal components.
@@
Finally, Figure 14 further shows the appropriateness of the seasonal ARTFIMA model by displaying a log-log plot of the periodogram and spectral density at the MAP. It is apparent that the seasonal ARTFIMA model fits the lower frequencies, $\text{log}(\omega)<-4$, and also models the seasonal components, shown in the zoomed-in section of the plot, as it captures the peaks and troughs of the periodogram accurately.
@@
Figure 14. Outer plot: Log frequency vs log power (log-log) plot of the periodogram and the spectral density function at the MAP for DLR with SARTFIMA$(2,d,\lambda,1)(0,0,1)_{48}$ errors. Inner plot: Same image but displaying moderate frequencies on a linear scale. The 95% credible interval of the spectral density estimated is the shaded blue region.
@@ 6. Conclusion and future research
We propose a dynamic linear regression model with an ARTFIMA error process that accounts for long-range dependence. We develop a frequency domain estimation method that utilises the computationally efficient FFT. The frequency domain method uses a likelihood based on asymptotic properties of the periodogram, and we demonstrate empirically that the resulting inference matches well with time domain likelihood methods. This is expected, as the case without exogenous variables has strong theoretical results indicating that the approximation error vanishes asymptotically.
We show empirically that the well-known phenomena of the spectral density of the ARFIMA model not fitting the periodogram data well for small frequencies carries over to the dynamic linear regression case. As such, we argue that dynamic linear regression models with ARFIMA errors are unsuitable for frequency domain methods, and their estimation needs time domain methods which are computationally expensive. In contrast, we show that the dynamic linear regression models with ARTFIMA errors fit the low-frequency spectrum well and are thus a computationally fast alternative to model long-range dependence.
Our proposed models demonstrate superior forecasting performance when the data exhibits complex long-range dependence and is on par with the ARMA error process for simpler data. This is expected as our model nests the dynamic linear regression model with ARMA errors. We do not see any signs of overfitting when using an ARTFIMA error process instead of an ARMA for the simpler dataset. In addition, we also observe that the posterior distribution of the regression parameter is significantly affected by not accounting for long-range dependence in the error process. In summary, we believe dynamic linear regression models with ARTFIMA errors are robust and should be routinely included in a forecaster’s toolkit.
Future research will extend our approach to modelling a multivariate response as in Villani et al., (2024). For small to moderate datasets, the Whittle log-likelihood is known to be severely biased. Exploring recent advances such as the debiased Whittle likelihood (Sykulski et al.,, 2019) for dynamic linear regression models can potentially enable our method to be used for small to moderately large datasets. Finally, the structure of our model can utilise subsampling MCMC methods (Quiroz et al.,, 2019, 2021) in the spectral domain (Salomone et al.,, 2020).